### v2

# HiRoPE: Length Extrapolation for Code Models Using Hierarchical Position

Abstract
Addressing the limitation of context length in large language models for code-related tasks is the primary focus of this paper. Existing LLMs are constrained by their pre-trained context lengths, leading to performance issues in handling long complex code sequences. Inspired by how human programmers navigate code, we introduce Hierarchical Rotary Position Embedding (HiRoPE), a novel approach that enhances the traditional rotary position embedding into a hierarchical format based on the hierarchical structure of source code. HiRoPE offers easy integration into existing LLMs without extra training costs. Our method is extensively evaluated with various LLMs, demonstrating stable performance in tasks such as language modeling and long code completion. We also introduce a new long code understanding task with real-world code projects, in hopes of promoting further development in this code-related field. Theoretically and experimentally, we find that HiRoPE also addresses the out-of-distribution issue in position encoding. Our HiRoPE significantly expands the context length capabilities of LLMs, enabling inference at lengths exponentially greater than the training length.
[…]
HiRoPE: Length Extrapolation for Code Models Using Hierarchical Position
Kechi Zhang,  Ge Li [^1] Corresponding authors.,  Huangzhao Zhang,  Zhi Jin [^1] Corresponding authors. Key Lab of High Confidence Software Technology (PKU), Ministry of Education School of Computer Science, Peking University, China {zhangkechi,lige,zhang_hz,zhijin}@pku.edu.cn
1 Introduction
Large language models (LLMs) such as LLaMA-2 Touvron et al. (2023b), and CodeLLaMA Rozière et al. (2023) have achieved significant performances in code-related tasks. These Transformer-based models excel in code comprehension and generation but face a notable challenge: the limitation of maximum context length. LLMs are typically pre-trained with a context length ranging from 2k to 16k tokens, which often proves insufficient for complex, extended source code. Exceeding this length limitation during inference may lead to performance degradation for these code models, particularly in tasks like project-level code completion or long code generation.
Various methods have been developed to extend the context window of LLMs. Some approaches involve fine-tuning on extensive texts Xiong et al. (2023); Chen et al. (2023b, a); Peng et al. (2023), which can be resource-intensive and potentially lead to overfitting and loss of performance on shorter sequences. There are also some training-free methods Xiao et al. (2023); Han et al. (2023); Ding et al. (2023). However, these methods usually use window attention rely on local information, and ignore the long dependency in code. It is essential to incorporate certain structural characteristics of the code into position encoding to efficiently model these long-distance code dependencies.
Figure 1: Illustration of the hierarchical position in source code, such as function-level and token-level positions. We also show a simplified abstract syntax tree of the code in the bottom left corner.
Our work diverges from these methods by focusing on the hierarchical information of source code in position encoding, inspired by how human programmers navigate code. Traditional positional encoding uses token counts for positioning, and treats code as plain text. However, human programmers often use hierarchical information in the code, representing positions in the code efficiently through multi-level hierarchical positions. We propose a hierarchical position approach that identifies token positions within specific levels, such as functions or statements. Figure 1 shows the comparison of the traditional position and our hierarchical position. It is clear that the hierarchical positional encoding, benefiting from the full utilization of structural information in the code, can more conveniently locate positional information within long code sequences. This method could more effectively model long dependencies in source code.
Following such inspirations, we introduce a novel approach, Hierarchical Rotary Position Embedding (HiRoPE), which enhances the popular rotary position embedding (RoPE) Su et al. (2024) into a hierarchical format. HiRoPE differentiates itself by extracting hierarchical information from the source code and splitting the RoPE dimension to represent different hierarchical levels. It simultaneously models token-level relative location and higher-level relative location information. We also add a window mechanism to ensure stability with short texts, aligning with traditional positional encoding.
HiRoPE is a plug-and-play solution, easily integrated into existing LLMs without additional training costs. Our extensive experiments with popular LLMs on tasks like language modeling and token completion in long code contexts CodeParrot (2022) demonstrate its effectiveness. We compare HiRoPE with existing length extrapolation methods using long code benchmarks such as LCC Guo et al. (2023) and RepoBench Liu et al. (2023a). We also introduce a new long code understanding task named code symbol understanding with real-world code libraries. Theoretically and experimentally, we find that HiRoPE effectively addresses the out-of-distribution issue Liu et al. (2023b) in position encoding. Our HiRoPE significantly expands the context length capabilities of LLMs, enabling inference at lengths exponentially greater than the training length. We believe our work with HiRoPE not only addresses a critical length limitation in LLM applications but also opens new avenues for long-structured data modeling research.
In summary, we make the following main contributions:
  • We propose Hierarchical RoPE (HiRoPE), enhancing the traditional rotary position embedding into a hierarchical format based on the hierarchical structure of source code, providing improved extrapolation capabilities.
  • We conducted comprehensive experiments with LLMs on various long code tasks involving language modeling and code completion. We also introduce a new long code understanding task with real-world code projects, in hopes of promoting further development in this code-related field.
  • We demonstrate that HiRoPE effectively addresses the out-of-distribution issue in position encoding, enabling inference at lengths exponentially greater than the training length.
2 Preliminary
We first introduce rotary position embedding in Transformer in Section 2.1. While existing work usually regards source code as plain text for modeling, we will also introduce the ignored hierarchical information in source code in Section 2.2.
2.1 Rotary Position Embedding in Transformer
Transformer models require explicit positional information to be injected, typically in the form of positional encodings, to represent the order of inputs. Recently, the rotary position embedding (RoPE) Su et al. (2024) has become one of the most popular and elegant position encoding strategies and is adopted by various LLMs Touvron et al. (2023a, b); Rozière et al. (2023). The main point of the RoPE method is using absolute position encoding to show relative position information. Formally, given a position index $m\in[0,L)$ and an embedding vector $\mathbf{x}:=[x_{0},x_{1},\ldots,x_{d-1}]^{\top}$, where $d$ is the dimension of the attention head, RoPE defines a complex function $\mathbf{f}(\mathbf{x},m)$ as follows:
$\begin{split}\mathbf{f}(\mathbf{x},m)=[(x_{0}+\mathrm{i}x_{1})e^{\mathrm{i}m\theta_{0}},(x_{2}+\mathrm{i}x_{3})e^{\mathrm{i}m\theta_{1}},\ldots,\\ (x_{d-2}+\mathrm{i}x_{d-1})e^{\mathrm{i}m\theta_{d/2-1}}]^{\top}\end{split}$ (1)
where $\mathrm{i}:=\sqrt{-1}$ is the imaginary unit and $\theta_{j}=10000^{-2j/d}$.
With RoPE, the self-attention score can be calculated as:
$\begin{split}&a(m,n)=\mathrm{Re}\langle\mathbf{f}(\mathbf{q},m),\mathbf{f}(\mathbf{k},n)\rangle\\ &=\mathrm{Re}\left[\sum_{j=0}^{d/2-1}(q_{2j}+\mathrm{i}q_{2j+1})(k_{2j}-\mathrm{i}k_{2j+1})e^{\mathrm{i}(m-n)\theta_{j}}\right]\\ &=\sum_{j=0}^{d/2-1}[(q_{2j}k_{2j}+q_{2j+1}k_{2j+1})\cos((m-n)\theta_{j})\\ &\quad+(q_{2j}k_{2j+1}-q_{2j+1}k_{2j})\sin((m-n)\theta_{j})]\\ &=:a(m-n)\end{split}$ (2)
Here $\mathrm{Re}\langle\dots\rangle$ denotes the real part function for a complex number. $\mathbf{q}$ and $\mathbf{k}$ are the query and key vector for a specific attention head. At each layer, RoPE is applied on both query and key embeddings for computing attention scores. We can observe that the calculated attention score is only dependent on relative position $m-n$ through trigonometric functions, which reflects the core of RoPE that uses the absolute position to represent the relative distance. Existing studies show that when dealing with long plain text, the RoPE will meet O.O.D issues where the value of $m-n$ during inference is unseen Liu et al. (2023b), leading to poor performances.
2.2 Hierarchical Position in Source Code
Most LLMs treat source code as plain text, processing it as if it were ordinary natural language. However, it is essential to take the structural information of code into mind. Source code can be transformed into abstract syntax trees, and these tree structures contain rich hierarchical position information. For example, the code snippets usually can be split into several class or function units, and each class/function contains various types of code blocks and statements. Figure 1 shows an illustration of the simplified abstract syntax tree for a code snippet in the bottom left corner. This higher-level positional information contains rich semantics of source code, making it easy for human programmers to locate and refer to different semantic parts. Therefore, in many existing program representation tasks, this hierarchical information plays a very important role Allamanis et al. (2018); Zhang et al. (2023). However, for today’s large language models, this high-level hierarchical positional information is almost ignored. In this paper, we try to incorporate this hierarchical information into the position encoding method.
3 Hierarchical RoPE
Figure 2: Overview of our HiRoPE. We transfer the existing position encoding method into a hierarchical format (i.e.,, function-level and token-level) and apply it across different dimensions. We also add a window mechanism to ensure performance stability (in this figure we set $L_{window}$ to 3).
In this paper, we propose HiRoPE, a hierarchical rotary position encoding for source code modeling. Our proposed HiRoPE requires two modified stages: ❶ Hierarchical Format: We first take the step to transfer the existing rotary position embedding into a hierarchical format. We verify our approach theoretically and find that the hierarchical format can bring stable extrapolation ability for RoPE. ❷ Window Mechanism. To ensure performance stability without further training, we also add a window mechanism, so that when dealing with short texts, our proposed method is consistent with the original positional encoding. An illustration of our HiRoPE is shown in Figure 2.
3.1 Hierarchical format
Unlike previous work that encodes the information of each position as a number index $m\in[0,L)$, we use a h-dimensional vector to represent the hierarchical position index from high-level to low-level, where $h$ is a hyperparameter that indicates how many levels of information we consider in the encoding. We begin with a simple case that we set $h=2$ and each token position can be represented as $(m_{1},m_{2})$. We use the higher and lower dimensions in RoPE respectively to represent these two hierarchical position indexes, so the Equation 1 can be rewritten as:
$\begin{split}&\mathbf{f}\textquoteright(\mathbf{x},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}m_{1}},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}m_{2}})=\\ &[(x_{0}+\mathrm{i}x_{1})e^{\mathrm{i}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}m_{1}}\theta_{0}},\ldots,(x_{d_{s}-2}+\mathrm{i}x_{d_{s}-1})e^{\mathrm{i}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}m_{1}}\theta_{d_{s}/2-1}},\\ &(x_{d_{s}}+\mathrm{i}x_{d_{s}+1})e^{\mathrm{i}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}m_{2}}\theta_{d_{s}/2}},\ldots,(x_{d-2}+\mathrm{i}x_{d-1})e^{\mathrm{i}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}m_{2}}\theta_{d/2-1}}]^{\top}\end{split}$ (3)
There are a total of $d$ dimensions in RoPE, and we use the lower $d_{s}$ dimensions to represent the hierarchical index of $m_{1}$, and the remaining dimensions to represent $m_{2}$. When we apply it to self-attention, we can get a new calculation of the attention score:
$\begin{split}&\mathrm{hierarchicalAttn}=\mathrm{Re}\langle\mathbf{f}(\mathbf{q},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}m_{1}},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}m_{2}}),\mathbf{f}(\mathbf{k},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}n_{1}},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}n_{2}})\rangle\\ &=\sum_{j=0}^{d_{s}/2-1}[(q_{2j}k_{2j}+q_{2j+1}k_{2j+1})\cos(({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}m_{1}-n_{1}})\theta_{j})\\ &\quad+(q_{2j}k_{2j+1}-q_{2j+1}k_{2j})\sin(({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}m_{1}-n_{1}})\theta_{j})]\\ &\quad+\sum_{j=d_{s}/2+1}^{d/2-1}[(q_{2j}k_{2j}+q_{2j+1}k_{2j+1})\cos(({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}m_{2}-n_{2}})\theta_{j})\\ &\quad+(q_{2j}k_{2j+1}-q_{2j+1}k_{2j})\sin(({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}m_{2}-n_{2}})\theta_{j})]\\ &=:a\textquoteright({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}m_{1}-n_{1}},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}m_{2}-n_{2}})\end{split}$ (4)
The attention score $a\textquoteright(\dots)$ ’ ( … ) we ultimately obtained through the inner product is quite elegant. It includes the relative position distance of various hierarchical levels, and this form can be similarly extended to the representation of more hierarchical positional structures, as shown in Equation 5. Specifically, when $h=1$, it remains the same as the original RoPE.
$\begin{split}&\mathrm{hierarchicalAttn_{h}}\\ &=\mathrm{Re}\langle\mathbf{f}(\mathbf{q},m_{1},\dots,m_{h}),\mathbf{f}(\mathbf{k},n_{1},\dots,n_{h}))\rangle\\ &=:a\textquoteright(m_{1}-m_{2},\dots,m_{h}-n_{h})\end{split}$ (5)
It indicates that the RoPE has the potential to be transformed into a hierarchical form, and we can use the hierarchical position index in the source code for this new form of RoPE, as shown in the left part of Figure 2.
The original RoPE sets $\theta_{j}$ to $10000^{-2j/d}$, which means that the lower the dimension, the higher its frequency, and the more emphasis on modeling the relative position information of short distances Pal et al. (2023). Therefore, in our hierarchical format, we use those low dimensions to represent token-level information, and high dimensions to represent higher-level hierarchical information, such as the function level or statement level in source code.
3.2 Window Mechanism
To ensure performance stability without further training, we follow existing extrapolation methods Xiao et al. (2023); Han et al. (2023) and add a window mechanism. Specifically, when dealing with short code snippets, we believe that existing LLMs have already mastered the ability to model these short semantic structures from vast pre-training code datasets. Therefore, when calculating the attention score, we directly use the original RoPE for those parts that are shorter than a specific length $L_{window}$. And for those long context parts that the distance is larger than $L_{window}$, we transfer them to the new hierarchical format by adding $L_{window}-1$ to each high-level distance. Our subsequent experiments have proved that even without any additional training, this window mechanism can bring strong stability, making the model’s performance applicable in various scenarios with arbitrary input lengths. An illustration of the final HiRoPE is shown in the right part of Figure 2.
4 Experiment Setup
Task Dataset Avg. Length Samples
Long Code Language Modeling CodeParrot 0-2048 1031.46 100
2048-4096 3667.76 100
4096-8192 7074.57 100
8192-16384 14353.94 100
Long Text Language Modeling ReRoPE-eval 21367.55 200
Code Symbol Understanding Real-world Code Project 12976.89 56
Long Code Completion LCC 17855.73 300
RepoBench 21103.42 300
Table 1: Statistics of the evaluation datasets
In this section, we aim to answer the following research questions through a series of experiments. Details of the evaluation dataset statistics are shown in Table 1.
RQ1. How is the language modeling capability of HiRoPE on long code sequences? We evaluate HiRoPE’s language modeling ability on CodeParrot-valid dataset CodeParrot (2022) in Section 5.1.
RQ2. How is the language modeling capability of HiRoPE on long natural language sequences? The natural language lacks the explicit hierarchical structure information found in code, so we have made some modifications: we set every 128 tokens as a segment, and encode it as higher-level position information. We use the evaluation dataset from ReRoPE-eval  [^1] https://github.com/bojone/rerope/blob/main/samples_15k.jsonl Su (2023) in Section 5.2. It is a dataset curated from Common Crawl Crawl (2023), refined by length-based selection criteria.
RQ3. How does HiRoPE perform in understanding real-world, long-code projects? To evaluate the effect of the method in real long-code scenarios, we design a new evaluation task on real code projects: Code Symbol Understanding in Section 5.3. Given a long code file, the model is required to output all the function names and class names defined in it. We extract long code files from popular open-sourced code repositories, especially those newly updated code projects to avoid data leakage. Details of these code projects are shown in Table 3.
RQ4. How does HiRoPE perform on existing benchmarks for long code completion? We further perform the evaluation using two long code completion benchmarks: LCC Guo et al. (2023) and RepoBench Liu et al. (2023a) in Section 5.4.
RQ5. What is the impact of various settings in HiRoPE? To demonstrate that each setting in the design of our HiRoPE works, we carry out extensive ablation studies that include the dimensions’ split settings, the window mechanism, and the high-level segment split strategy in Section 5.5.
4.1 Base LLMs
LLaMA-2 ShearedLLaMA TinyLLaMA Vicuna
Para. 7B 1.3B 1.1B 7B
$L_{pretrain}$ 4096 4096 2048 2048
Vocab Size 32000 32000 32000 32000
Hidden Size 4096 2048 2048 4096
Attention Head 32 16 32 32
RoPE Dim 128 128 64 128
Table 2: Statistics of base LLMs
The models used include LLaMA-2 (7B) Touvron et al. (2023b), Sheared-LLaMA (1.3B) Xia et al. (2023), TinyLlama (1.1B) Zhang et al. (2024), and Vicuna (7B) Chiang et al. (2023). This model choice is driven by their widespread use and popularity, as well as the constraints of our computing capabilities. Details are provided in Table 2.
4.2 Baselines
Considering training on long context sequences is resource-intensive and time-consuming, we focus on those popular length extrapolation methods without training, including NTK bloc97 (2023), ReRoPE Su (2023) and Self-Extend Jin et al. (2024). ReRoPE sets a window when calculating the relative distance, and clips those distances outside the window. Self-Extend uses group attention with the floor operation to calculate relative distances. These methods have shown impressive performance on long context language modeling. We also make comparisons with the original RoPE method, which is denoted as “origin” in the first row of Table 4, 5 and 6.
4.3 Inference Settings
In our experiments, our HiRoPE uses a two-layer hierarchy, accounting for the position index at the token and function/class levels of the source code based on tree-sitter Brunsfeld et al. (2024). For long context in natural language, we make some modifications and set every 128 tokens as a higher-level segment. We set the split dimension half of the total: $d_{s}=0.5*d_{total}$, and choose a window length: $L_{window}=512$. We keep the hyperparameters the same for those state-of-the-art baselines for a fair comparison. We use greedy search decoding for generation. We use 4 A6000 GPUs for all experiments.
4.4 Details of Code Symbol Understanding task
In RQ3, we design a new task: Code Symbol Understanding. Given a long code context, the model is required to output all the function names and class names defined in it. We extract long code files from popular open-sourced code repositories, especially those newly updated code projects to avoid data leakage. We construct a static analysis tool to get the abstract syntax tree of each code file, and then get all defined function and class names in it as the ground-truth output symbols. Details of these code projects are shown in Table 3. We show the number of files we extracted from each project as well as the average length of these code files. We also show the average number of symbols and their location statistics in the table.
In Figure 3 we show an illustration of our proposed new task. We replace the input_code part with each code file, and use the task prompt to guide models to extract and output all defined function and class names in input code. We also show an example output in the figure.
Github Repo File Nums File Length Avg. Symbols Min. Symbol Loc Max. Symbol Loc
ddbourgin/numpy-ml 2 14055 15 927 10845
gradio-app/gradio 1 12938 11 272 12265
huggingface/accelerate 4 12814.5 11.5 379 12500
huggingface/diffusers 4 13324.75 16.2 276 8347
huggingface/optimum 1 13491 11 580 12868
huggingface/peft 1 15457 11 514 8525
huggingface/transformers 18 12919.4 12.7 239 14514
langchain-ai/langchain/ 2 14195 11.5 285 10790
numpy/numpy 6 11158.8 14 100 12026
tensorflow/tensorflow 17 13191.7 13.4 285 14920
Total 56 12976.89 13.17 100 14920
Table 3: Statistics of Code Symbol Understanding task. We show the number of files we extracted from each project as well as the average length of these code files. We also show the average number of symbols and their location statistics in the table.
Figure 3: Illustration of Code Symbol Understanding task. We use the task prompt to guide models to extract and output all defined function and class names in input code.
5 Results and Analyses
5.1 Long Code Language Modeling
Dataset Length: 0-2048 2048-4096 4096-8192 8192-16384
Para. $L_{pretrain}$ loss $\downarrow$ ppl $\downarrow$ acc $\uparrow$ loss $\downarrow$ ppl $\downarrow$ acc $\uparrow$ loss $\downarrow$ ppl $\downarrow$ acc $\uparrow$ loss $\downarrow$ ppl $\downarrow$ acc $\uparrow$
LLaMA-2 7B 4096 origin 0.8579 2.3583 0.8065 0.9820 2.6698 0.7551 nan nan nan nan nan nan
NTK 1.1107 3.0365 0.7472 1.0469 2.8488 0.7414 0.9858 2.6800 0.7729 1.0181 2.7678 0.7630
ReRoPE 0.8593 2.3615 0.8054 0.7278 2.0705 0.8121 0.6633 1.9411 0.8411 0.7187 2.0518 0.8243
Self-Extend 0.8588 2.3604 0.8055 0.7209 2.0562 0.8147 0.6519 1.9192 0.8441 0.6983 2.0103 0.8290
HiRoPE 0.8586 2.3598 0.8060 0.7185 2.0514 0.8153 0.6482 1.9121 0.8452 0.6821 1.9780 0.8332
ShearedLLaMA 1.3B 4096 origin 1.2874 3.6235 0.7341 1.3103 3.7074 0.7019 3.8381 46.4375 0.4866 5.6958 297.6160 0.3211
NTK 1.6242 5.0744 0.6607 1.5047 4.5029 0.6619 1.3708 3.9383 0.7015 1.3657 3.9183 0.6964
ReRoPE 1.2897 3.6316 0.7332 1.0497 2.8567 0.7560 0.9699 2.6376 0.7846 1.0044 2.7303 0.7716
Self-Extend 1.2892 3.6300 0.7337 1.0428 2.8371 0.7586 0.9568 2.6034 0.7874 0.9804 2.6656 0.7768
HiRoPE 1.2888 3.6285 0.7338 1.0382 2.8242 0.7600 0.9514 2.5894 0.7885 0.9660 2.6273 0.7811
TinyLlama 1.1B 2048 origin 1.0788 2.9410 0.7594 4.1732 64.9235 0.4506 6.6603 780.8009 0.2630 7.9938 2962.5881 0.1682
NTK 1.1837 3.2664 0.7405 1.0952 2.9899 0.7256 0.9719 2.6430 0.7733 1.0021 2.7240 0.7626
ReRoPE 0.9703 2.6388 0.7877 0.8251 2.2821 0.7905 0.7685 2.1565 0.8210 0.8275 2.2877 0.8040
Self-Extend 0.9698 2.6375 0.7856 0.8123 2.2530 0.7931 0.7577 2.1333 0.8235 0.8119 2.2521 0.8073
HiRoPE 0.9743 2.6493 0.7881 0.8268 2.2861 0.7981 0.7683 2.1562 0.8208 0.8040 2.2345 0.8094
Vicuna 7B 2048 origin 1.1787 3.2502 0.7551 4.6046 99.9449 0.4473 7.7207 2254.4730 0.2597 9.9449 20846.3927 0.1601
NTK 1.3417 3.8255 0.7150 1.2587 3.5208 0.7068 1.1809 3.2573 0.7344 1.1912 3.2911 0.7285
ReRoPE 1.0716 2.9201 0.7800 0.8760 2.4012 0.7912 0.8138 2.2566 0.8182 0.8580 2.3585 0.8023
Self-Extend 1.0710 2.9183 0.7802 0.8735 2.3953 0.7891 0.7988 2.2228 0.8220 0.8351 2.3049 0.8066
HiRoPE 1.0707 2.9174 0.7799 0.8724 2.3926 0.7903 0.8002 2.2261 0.8213 0.8314 2.2965 0.8080
Table 4: Language Modeling Ability on CodeParrot-valid dataset. “nan” indicates that the model performs significantly poor on the given setting.
Language modeling is the most fundamental and the least requirement for a LLM. We evaluate HiRoPE’s language modeling ability on CodeParrot-valid dataset. We divide the original dataset into different length intervals. The experiment results are shown in Table 4. A smaller loss and a smaller ppl indicate a stronger language modeling capacity of the corresponding model. Conversely, a larger acc suggests a stronger capability of the respective model in code completion on the given dataset. The origin indicates that we directly use the original setting of the model to evaluate.
Experiments show that original LLMs perform badly on the long code language modeling task. Even when the length slightly exceeds the pre-training length, the ppl of all models exceed 45, demonstrating their essential lack of modeling and understanding capabilities for longer codes. When we apply length extrapolation methods, all methods can reduce the loss and ppl into an acceptable range for those long source codes. Specifically, for ultra-long code sequences (length over 8192), HiRoPE achieves the best results in all metrics and under all settings. This fully demonstrates the advantages of our HiRoPE in modeling long sequence codes. Our method also shows generalization abilities. The four models evaluated have differences in model parameters, pre-training data, and pre-training length, yet our method has shown very good results on all these models.
It is worth noting that our method does not impair the model’s performance on shorter code. We noted that some popular length extension methods, such as NTK, can impair the performance on short code. Thanks to our window mechanism, our method stays on par with the baseline model on shorter datasets (length 0-2048) and even surpasses the baseline on some metrics. HiRoPE demonstrates consistently excellent language modeling capabilities in various code length scenarios.
5.2 Long Text Language Modeling
last loss $\downarrow$ last ppl $\downarrow$ last acc $\uparrow$ all loss $\downarrow$ all ppl $\downarrow$ all acc $\uparrow$
TinyLlama origin 9.3083 >1000 0.0194 7.3486 >1000 0.1466
NTK 2.1338 8.4468 0.5553 2.2745 9.7234 0.53
ReRoPE 1.8448 6.327 0.6008 1.903 6.7062 0.5867
Self-Extend 1.7904 5.9916 0.6089 1.8749 6.5201 0.5908
HiRoPE 1.7829 5.9474 0.6102 1.8717 6.4991 0.5913
ShearedLLaMA origin 8.5027 >1000 0.0412 5.5237 250.5 0.274
NTK 2.3596 10.5863 0.5159 2.4056 11.0853 0.5059
ReRoPE 1.7952 6.0205 0.605 1.8514 6.3687 0.5932
Self-Extend 1.7662 5.8486 0.6105 1.8348 6.264 0.5966
HiRoPE 1.7622 5.8252 0.6113 1.8332 6.2536 0.5963
Table 5: Language Modeling Ability on ReRoPE-eval dataset. In addition to calculating metrics on all tokens (refer to "all_…"), we also record metrics on the last 2048 tokens of each data (refer to "last_…" ).
In addition to testing the ability of language modeling on long code, we also evaluate its effects on long natural language texts. We set every 128 tokens as a high-level segment. We use the ReRoPE-eval dataset, and results are shown in Table 5. In addition to calculating metrics on all tokens (refer to "all_…" in the table), we also record metrics on the last 2048 tokens of each data (refer to "last_…").
We find that our HiRoPE can also achieve significant improvement. HiRoPE achieves the best results on almost all metrics. Another interesting observation is that, given sufficient context, the model can utilize this contextual information to perform better when generating later tokens. That is to say, the metrics of "last_…" should be better than those of "all_…". However, we observe that for the original model, the situation is contrary to this. We attribute this to the fact that the original model’s ability to model long sequence languages is so poor that it can’t utilize that distant contextual information at all. Our HiRoPE can significantly improve the model’s ability to handle long codes and textual data, without requiring any training, reflecting its practical value.
5.3 Code Symbol Understanding
Code Symbol Understanding (Recall $\uparrow$) Long Code Completion (Edit Sim $\uparrow$)
LCC RepoBench
0-4k 4k-8k >8k 0-4k 4k-8k >8k
LLaMA-2 origin 0.0012 54.5 4.36 4.08 8.29 6.79 6.59
ReRoPE 0.0837 65.83 67.43 63.22 52.82 47.85 45.38
HiRoPE 0.0911 66.61 69.93 65.38 52.20 53.30 51.24
ShearedLLaMA origin 0.0067 27.17 3.33 2.52 4.39 2.94 2.35
ReRoPE 0.0743 35.56 36.1 36.91 34.03 37.37 33.44
HiRoPE 0.0809 46.13 51.67 46.33 40.17 39.98 39.52
TinyLLaMA origin 0.0067 17.29 5.45 6.28 7.91 7.53 7.07
ReRoPE 0.1214 49.22 57.20 53.11 37.72 40.50 39.38
HiRoPE 0.1415 35.17 42.83 49.92 42.48 43.53 39.82
Vicuna origin 0.0067 18.47 2.56 2.76 3.67 2.49 2.32
ReRoPE 0.0636 57.95 59.73 58.52 42.78 43.65 45.23
HiRoPE 0.0721 63.42 62.01 64.42 37.10 42.30 45.93
Table 6: Performance on Code Symbol Understanding and Long Code Completion.
To evaluate the effect of the method in real long-code scenarios, we have designed an evaluation task for real code projects: Code Symbol Understanding. Given a long code context, the model is required to output all the function names and class names defined in it. These pre-defined functions and classes are reused frequently in actual code development. For code models, understanding which functions and classes are defined in the code project is a basic capability requirement. This task is inspired by the popular "Needle in a Haystack" synthetic evaluation gkamradt (2023), but our code symbol understanding task is more realistic and code-related. Task examples are shown in Figure 3. We use recall as the evaluation metric.
Our experiments in Table 6 have proven that this seemingly simple task is extremely difficult for LLMs. We also evaluate this task using GPT-3.5-16k GPT-3.5 (2023) and find that its result is only 0.72. All these LLMs are not ideal in this more realistic code symbol understanding task. Our HiRoPE has been improved from the perspective of positional encoding, enabling the model to perceive structural hierarchy changes in the code, thus achieving relatively good results. Compared to the original models, our HiRoPE can achieve almost a hundredfold improvement on average across four models. We release our dataset in hopes of promoting further development in this code-related field.
5.4 Long Code Completion
We further perform the evaluation using two real-world long code completion benchmarks: LCC and RepoBench. Given a long code context, the model is required to generate the complete next line of code. We follow the experiment settings in Longbench-E Bai et al. (2023) and use edit similarity as metrics. Results are shown in Table 6.
Our HiRoPE also achieves stable improvements on this long code-related task. The input code context is filled with various predefined functions and classes. Our method can effectively sense these contents, handle these complex dependencies, and successfully use those functions that are defined far away during generation. Under two dataset scenarios, our method consistently outperforms other baseline settings. The results reflect the practicality and generalization ability of our method.
5.5 Ablation Study
Figure 4: Ablation Studies including the settings of the dimension split, the window mechanism, and the high-level segment split strategy.
In our experiments, ❶ we set the split dimension half of the total $d_{s}=0.5*d_{total}$ and ❶ choose a window length ❷ $L_{window}=512$. ❸ Our hierarchical position includes both the token and function/class levels of the source code. We further carry out extensive ablation studies including these settings and results are shown in Figure 4.
We choose different dimension splitting ratios to observe the performance of LLaMA-2 and TinyLLaMA in terms of ppl on CodeParrot [4k-8k]. When the split ratio is 1, the HiRoPE degenerates into the original model. Specifically, both models show significant fluctuations in ppl between ratios of $[0.6,0.7]$. We will explore the reason in Section 6.1.
We change the window length to observe the performance of LLaMA-2 on CodeParrot [0-2k] and [4k-8k]. For shorter code data, we observe a decreasing trend in ppl as the window size increases. This validates that the window mechanism can allow the model to retain its original computational mechanisms and better handle short-distance dependencies. For longer code data, ppl behaves anomalously when the window size is very small. This also indicates that window mechanisms play a key role in modeling long-distance dependencies.
We change the high-level segment split strategy. In addition to dividing the hierarchy at the func-level, we also try to split at the code statement level as well as implementing a strategy of splitting continuous n-tokens as a high-level segment (n = 128, 512, 1024). Experiments show that dividing at the function level achieves the best results. The semantics within a function are relatively similar, while the semantics between functions usually vary greatly. It is necessary to divide long code sequences into levels according to functions and classes.
6 Discussion
6.1 Mitigating Out-of-Domain Issues in Long Contexts
Existing work Liu et al. (2023b) shows that LLMs fail on input out of the pretraining context window because of the Out-Of-Distribution (O.O.D) issues. We take the inspiration to explain it from a cyclical perspective. In RoPE, each dimension is composed of trigonometric functions, and its period can be denoted as $T_{j}=\frac{2\pi}{\theta_{j}}=2\pi*10000^{\frac{2j}{d}}$. Only those dimensions that have been completely trained within the pre-training length can be considered as a reliable part for extrapolation, others would encounter O.O.D issues when dealing with problems of extrapolation. We can then get those reliable dimensions by calculating $T_{j}<L_{pretrain}$. The calculated dim split is 0.70 and 0.63 for the four models in our experiments (Table 7). We are surprised to find that it is similar to the ratio we obtained in the ablation study in Figure 4. Our HiRoPE uses those high dimensions to represent higher-level position index information, and properly applies them to smaller input numbers, thus mitigating O.O.D issues in long code contexts.
The traditional RoPE uses a number $m$ as the position index to represent position information. Due to the O.O.D problem in high dimensions, its reliable range is $\{m\in[0,L_{pretrain}]\}$. In our HiRoPE, we use a two-layer hierarchy as $(m_{1},m_{2})$ and the reliable range is $\{m_{1}\in[0,L_{pretrain}],m_{2}\in[0,L_{pretrain}]\}$. It proves that under ideal circumstances, HiRoPE can effectively extrapolate to the length of $L^{h}$ in an exponential ratio, where $h$ is the hierarchy layer. We attempt to explore the upper limit of our HiRoPE’s extrapolation performance in the experiment in the next Section 6.2.
6.2 Upper Limit of HiRoPE’s Performance
Figure 5: Performance of Short-ShearedLLaMA on CodeParrot dataset. The training length is set to 128. The results suggest our method has the potential to extrapolate code models at an exponential length.
Due to computational resource constraints, we made the following modifications based on Section 5.1: ❶ Firstly, in order to obtain a base model of suitable length, we designed a training strategy to obtain a shorter context length LLM, named ShortLLM: We use the position interpolation Chen et al. (2023b) method reversely to reduce the input length of some mainstream models to $L_{short}=128$ at a smaller training cost. Specifically, given a position index $m$ in the original RoPE, we use the new index $\alpha_{short}*m$ to replace it as shown in Table 8. We fine-tune short models for 1000 steps. ❷ We resample the CodeParrot-valid dataset, further refining it into smaller distance ranges, each range containing up to 50 test samples. The results are shown in Figure 5 and 6.
Our trained ShortLLM successfully demonstrates the expected performance: the performance drastically decreases after surpassing the training length $L_{short}=128$. We then apply our HiRoPE as well as the baseline ReRoPE. Our HiRoPE demonstrates a more stable trend on long code sequences, even at the position close to $L_{short}^{2}$. It suggests our method has the potential to extrapolate code models at an exponential length.
7 Related Work
Existing large language models are originally trained with fixed context sizes. When dealing with longer sequences, the model’s performance may decrease quite drastically. Recent studies have explored ways to expand the context length. Position Interpolation Chen et al. (2023b) linearly down-scales the input position indices to match the original context window size of LLMs with several training steps. Similar studies Chen et al. (2023a); Peng et al. (2023); Chen et al. (2023c); Guo et al. (2023) also require fine-tuning. However, these methods need additional tuning in longer contexts or face a disastrous collapse after the extrapolation bound. There are also some approaches without training. Some studies use window attention to clip the long sequences such as Xiao et al. (2023); Han et al. (2023); Ding et al. (2023)). However, these methods rely on local information and may not effectively expand the context window, struggling with long dependencies in code. Recently, some methods have explored modifying the relative distance to extend the extrapolation length bloc97 (2023); Su (2023); Jin et al. (2024) and focus on the natural language text. We pursue the research line of training-free methods and propose considering the structural information of the code when modeling the position. Our work is inspired by the hierarchical embedding in our former work Zhang et al. (2023). We notice hierarchical positions are used in some concurrent work He et al. (2024); Lu et al. (2024) in NLP and CV domains. We expand the traditional RoPE method into a hierarchical format and prove the effectiveness of our HiRoPE through theoretical derivations and practical experiments.
8 Conclusion
We propose HiRoPE, a training-free solution to the context length limitation in LLMs for long code modeling. We integrate the hierarchical structure of source code into position encoding of LLMs. Experiments demonstrate that HiRoPE achieves stable improvements on diverse code-related tasks. Our work not only addresses a critical limitation in LLM applications but also opens new avenues for long structured data modeling research.
9 Acknowledgments
This research is supported by the National Key R&D Program under Grant No. 2023YFB4503801, the National Natural Science Foundation of China under Grant No. 62072007, 62192733, 61832009, 62192730, and the Major Program (JD) of Hubei Province (No.2023BAA024).
Limitations
There are several limitations to our work that we aim to address:
Firstly, constrained by computational resources, we choose models below 7B for experiments. The four models evaluated have differences in model parameters, pre-training data, and pre-training length, yet our method has shown very good results on all these models. We will attempt to conduct experiments on models with larger parameters and more complex structures to promote the development of the LLM community.
Next, our discussion on the upper limit of the HiRoPE’s performance tends to lean towards theoretical derivation. We have designed a set of ShortLLM experiments to prove our conclusions. It suggests our method has the potential to extrapolate code models at an exponential length, so for the LLaMA-2 model with $L_{pretrain}=4096$, we can theoretically extrapolate its length to $L_{pretrain}^{2}\approx 16,000,000$. We are not clear whether some settings will implicitly affect the performance of the model. We will continue to explore the robustness of this experimental idea and try to explore the maximum performance of our method on real LLMs.
Appendix A Details of ShortLLM experiments
Figure 6: Performance of ShortLLMs on CodeParrot dataset. The training length is set to 128.
A.1 Theoretical calculation for O.O.D issues
According to the analysis in Section 6.1, we can calculate the reliable dimensions for each model as:
$\begin{split}Split\%=log_{10000}\frac{L_{pretrain}}{2\pi}\end{split}$ (6)
$L_{pretrain}$ RoPE Dim Split % Split Dim
LLaMA-2 4096 128 0.70 90.05
ShearedLLaMA 4096 128 0.70 90.05
TinyLLaMA 2048 64 0.63 40.21
Vicuna 2048 128 0.63 80.42
Table 7: Theoretical calculation of the reliable extrapolation dimension.
The theoretical calculation results are shown in Table 7.
A.2 ShortLLM Training
Para. $L_{pretrain}$ $L_{short}$ $\alpha_{short}$
LLaMA-2 7B 4096 128 32
ShearedLLaMA 1.3B 4096 128 32
TinyLLaMA 1.1B 2048 128 16
Vicuna 7B 2048 128 16
Table 8: Model Statistics for ShortLLM experiments.
In order to obtain the ShortLLM of suitable length, we use the position interpolation Chen et al. (2023b) method reversely. Given a position index $m$ in the original RoPE, we use the new index $\alpha_{short}*m$ to replace it as shown in Table 8. We sample the CodeParrot-train dataset and filter the data length less than $L_{short}$. We set the global batch size to 64 and fine-tune the models for 1000 steps.
A.3 ShortLLM Performances
After training, we apply our HiRoPE to these ShortLLM models. Figure 6 shows the experiment results of the ShortLLM. We can observe that for all models, Our HiRoPE can maintain good stability as the length of the input code increases. Even if the baseline ReRoPE method gradually becomes unstable under some experimental settings (such as on Short-TinyLLaMA), our method can resist these performance declines.